If (3x+5y)/(2x-y)=(12x-y)/(5x-3y) then find x:y
Answer by stanbon(75887)

If (3x+5y)/(2x-y)=(12x-y)/(5x-3y) then find x:y
--------------
Cross-multiply:
= (3x+5y)(5x-3y) = (2x-y)(12x-y)
---
= 15x^2 + 6xy - 15y^2 = 24x^2 - 14xy + y^2
-----
= 9x^2 - 20xy +16y^2
---
Not factorable in Real Number System
==========================================
Cheers,
Stan H.